4) Find the inverse Laplace transform of 5/(s-1)(s+4) without using partial fractions.


1
Expert's answer
2021-09-28T00:58:21-0400

"=((\frac{s}{5}−\frac{1}{5})(s+4))"

Its laplace form is;


"=\frac{-4s^2+3s+2}{5s^3} \space answer"